Find the electric field midway between a 2 -μC point charge which are 2.5 m apart.

1
Expert's answer
2022-02-17T04:36:30-0500

Given:

"q=2\:\rm \mu C=2*10^{-9}\: C"

"r=1.25\:\rm m"


The electric field of a point charge

"E=k\frac{q}{r^2}=9*10^9*\frac{2*10^{-9}}{1.25^2}=11.5\:\rm V/m"